75.0ml of 1.00M HCl is added to 100ml of 1.00M NaOH. What is the pH of the resulting solution? (Please show your work.)
1
Expert's answer
2020-10-14T06:02:41-0400

75.0ml of 1.00M HCl is added to 100ml of 1.00M NaOH


mols of HCl present = 75 ml * 1 M = 75 milli mols


mols of NaOH present = 100 ml * 1 M = 100 milli mols


Since both acid and base are strong, and they react in 1:1 mol ratio,


75 milli mols of HCl will react with 75 milli mols of NaOH


Thus there will be , 100 - 75 millimol = 25 millimol NaOH remains in the

solution. 

Thus [NaOH] = 25 m mol/(100 +75 ml) = 0.14285 M


NaOH--> Na + + OH-


thus [OH-] = 0.14285 M

pOH = -log [OH-] = 0.845

pH = 14-pOH = 13.15


ANSWER

pH = 13.2